Martinez Corporation has fixed costs of $2,859,600. It has a unit selling price of $8.00, unit variable costs of $4.40, and a target net income of $1,590,000. Calcuate the required sales in units to achieve its target net income.

Required sales= ___ units
Answer #1

Contribution margin=Sales-Variable cost

=8-4.4=$3.6 per unit

Target Contribution margin=Fixed cost+Target profits

=2,859,600+1,590,000=$4449600

Hence required sales=4449600/3.6

=1,236,000 units